OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos 2(theta) = cos^2(theta) - sin^2(theta)

OpenStudy (raden):

cos(2θ) = cos(θ+θ) = cosθcosθ - sinθsinθ = cos^2 θ - sin^2 θ hoped that make sense

OpenStudy (anonymous):

thanks!